OpenStudy (anonymous):

hello need help,plotting differentiattion graph,but am totally confuse in this case as my cn's are {-2,0) and my hyper cn's {-3,0).how do i do this in the graph?

OpenStudy (anonymous):

This are my given equation\[f(x)=2(x+1)/x^2 , f \prime(x)=-2(x+2)/x^3 and f \prime \prime(x)=4(x+3)/x^4\]

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

i'm confused. what are you trying to do?

OpenStudy (anonymous):

derivatives are correct...

OpenStudy (anonymous):

Application of differentiation Graphing

OpenStudy (anonymous):

oh, I see. graph f using f' and f''. is that right?

OpenStudy (anonymous):

yeah exactly!!!

OpenStudy (anonymous):

have taken some steps already,can given u my answers if u like,but cant actually do the graphing cos my asnwers look confusing to me

OpenStudy (anonymous):

ok let's start with f first. domain of f is any x except x=0... you'll have an asymptote there.

OpenStudy (anonymous):

are there any zeros of f?

OpenStudy (anonymous):

yeah positive in both left and right

OpenStudy (anonymous):

my horizontal asympto is 0 y=0

OpenStudy (anonymous):

correct for horizontal asymptote. but do you have any zero's? (yes, horizontal asymptotes CAN be crossed)

OpenStudy (anonymous):

as in? u mean my intercepts? or my relative(local)extrema?

OpenStudy (anonymous):

zeros = x-intercept(s)

OpenStudy (anonymous):

(-1,0)

OpenStudy (anonymous):

good. this is what we have so far...